OpenStudy (anonymous):

if the same iron cube weighs 7.9kg, what is its density in g/cm3

OpenStudy (aaronq):

\(\sf density=\dfrac{mass}{volume}\) you have to convert kg to g; 1 kg = 1000g

OpenStudy (anonymous):

what is the density of a cube of water measuring 2cm * 4cm * 1cm, with a mass of 8g ?

OpenStudy (aaronq):

you find the volume by multiplying the sides together. then divide the mass by the volume

OpenStudy (anonymous):

density = \[mass \div volume\] so volume of cube is \[side \times side \times side\] and the dimension you have specified are for cuboid